OpenStudy (anonymous):

how do you get the other sides of a right triangle given the hypotenuse? is it by Pythagorean theorem?

OpenStudy (anonymous):

to find the hypotenuse it is a^2 +b^2= c (c is the hypotenuse) so i would imagine you would do the inverse of that

OpenStudy (anonymous):

for a right triangle if you are given an angle (other than the 90) you can use trig. With x=the angle measure sin(x)=(opposite side)/(hypotenuse) and cos(x)=(adjacent side)/(hypotenuse) solve for the side you are looking for and you should get your answer.

OpenStudy (anonymous):

^She's right. There are infinite possibilities for the other two sides for a given hypotenuse. But if the other two sides are given, only one hypotenuse can exist.

OpenStudy (anonymous):

The Pythagorean theorem tells you this: \[C² = A² + B²\]Let C be the hypontenuse then and let's say you want to get the side A then: \[A² = C² - B²\]Thus: \[A = \sqrt{C² - B²}\]That's when you are given C and B. Otherwise, with angles, use what TMane said.